# Coupled Lindblad pseudomode theory for simulating open quantum systems

### v1

@@ Abstract
Finally, we compute the absorption spectra using the coupled Lindblad pseudomode theory for a dimer model with three states, $\mathinner{|{g}\rangle},\mathinner{|{\epsilon_{1}}\rangle},$ and $\mathinner{|{\epsilon_{2}}\rangle}$, and $\hat{H}_{\text{S}}=\sum_{i=1}^{2}\epsilon_{i}\mathinner{|{\epsilon_{i}}\rangle}\!\mathinner{\langle{\epsilon_{i}}|}+J(\mathinner{|{\epsilon_{1}}\rangle}\!\mathinner{\langle{\epsilon_{2}}|}+\mathinner{|{\epsilon_{2}}\rangle}\!\mathinner{\langle{\epsilon_{1}}|})$. The environment is coupled independently to each excited state via the operators, $\hat{S}_{i}=\mathinner{|{\epsilon_{i}}\rangle}\!\mathinner{\langle{\epsilon_{i}}|}$, using the parameters of Ref. [9]. We consider two different spectral densities: $J_{0}(\omega)$, a broad spectrum, and $J_{1}(\omega)$, which features an additional sharp peak over $J_{0}(\omega)$ (details in Supplemental Material (SM) [53]). We compute the absorption spectrum, $S(\omega)=\omega\operatorname{Im}\left(\int_{0}^{\infty}\mathrm{i}C_{\hat{\mu}}(t)\mathrm{e}^{\mathrm{i}\omega t}\mathrm{d}t\right)$, derived from the dipole-dipole correlation function $C_{\hat{\mu}}(t)$ with $\hat{\mu}=\sum_{i}\mathinner{|{\epsilon_{i}}\rangle}\!\mathinner{\langle{g}|}+\mathinner{|{g}\rangle}\!\mathinner{\langle{\epsilon_{i}}|}$ and an initial state $\hat{\rho}_{\text{S}}(0)=\mathinner{|{g}\rangle}\!\mathinner{\langle{g}|}$. The result is shown in Fig. 3 compared to results from the Lorentzian pseudomode theory. Notably, the absorption spectrum exhibits a sharp peak in the negative frequency region at zero temperature, which becomes narrower and sharper as $N$ increases, highlighting the convergence behavior of our approach. Accurately capturing the presence of sharp resonances and broad features in the spectrum requires a faithful reconstruction of the spectral density, which our method achieves effectively. In contrast, the Lorentzian pseudomode fails to reproduce the broad component and misplaces the positions of several spectral peaks.
@@
Conclusions and outlook.– We argue that the coupled Lindblad pseudomode framework possesses all the desirable features of a well-designed pseudomode theory. The dynamics can be realized as a quantum channel, making them inherently stable. In both theory and practice, only a small number of pseudomodes is needed to achieve accurate results with $\text{polylog}(T/\varepsilon)$ scaling. Moreover, the pseudomodes can be constructed via a robust algorithm. We illustrate the method using the spin-boson model as an example, but it is equally applicable to fermionic environments (see SM [53]). We anticipate that with the improved understanding and robustness provided by the techniques developed in this work, the coupled Lindblad framework will prove broadly useful across a wide range of open quantum system applications. These include the simulation of ultra-strong coupling regimes [54], the use of pseudomodes as an impurity solver in dynamical mean-field theory [55, 56], and the simulation of condensed-phase chemical dynamics on quantum platforms such as trapped-ion devices [25, 26].
@@
(Dated: June 12, 2025)
I Coupled Lindblad pseudomodes theory for multi-site cases
@@
Therefore the coupled Lindblad dynamics takes the following form:
$\frac{\mathrm{d}\hat{\rho}}{\mathrm{d}t}=-\mathrm{i}[\hat{H}_{\text{S}}+\hat{H}_{\text{A}}+\hat{H}_{\text{SA}},\hat{\rho}]+\boldsymbol{D}_{\text{A}}(\hat{\rho}),$ (S3)
@@
II Coupled Lindblad pseudomode theory for fermionic systems
@@
[…]
@@
Here $J(\omega)=f_{i}(\omega)\overline{f_{j}(\omega)}$ is the bath spectral density, $\mu$ is the chemical potential, $\beta$ is the inverse temperature, and $f_{\text{FD}}^{\mu,\beta}(\omega)=\frac{1}{1+\mathrm{e}^{\beta(\omega-\mu)}}$ is the Fermi-Dirac function. To apply the pseudomode theory for fermionic cases, the key is to account for both lesser and greater real-time hybridization functions arising in the problem. This is similar to what is done in quasi-Lindblad theories [7, Sec II. D]. We conduct fitting for both $\Delta^{<}(t)$ and $\Delta^{>}(t)$ as follows:
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}_{\text{SA}}$ $=-\mathrm{i}[\hat{H}_{\text{S}},\hat{\rho}]-\mathrm{i}[\hat{H}_{\text{A}_{1}}+\hat{H}_{\text{SA}_{1}},\hat{\rho}]-\mathrm{i}[\hat{H}_{\text{A}_{2}}+\hat{H}_{\text{SA}_{2}},\hat{\rho}]+\boldsymbol{D}_{\text{A}_{1}}(\hat{\rho})+\boldsymbol{D}_{\text{A}_{2}}(\hat{\rho}),\quad$ (S7)
@@
$\boldsymbol{D}_{\text{A}_{1}}(\hat{\rho})=\sum_{k,l=1}^{N_{2}}\Gamma^{<}_{kl}(2\hat{c}_{k}^{\dagger}\hat{\rho}\hat{c}_{l}-\{\hat{c}_{l}\hat{c}_{k}^{\dagger},\hat{\rho}\}),\quad\boldsymbol{D}_{\text{A}_{2}}(\hat{\rho})=\sum_{k,l=1}^{N_{2}}\Gamma^{>}_{kl}(2\hat{d}_{l}\hat{\rho}\hat{d}_{k}^{\dagger}-\{\hat{d}_{k}^{\dagger}\hat{d}_{l},\hat{\rho}\}).$
@@
III TD-DMRG based simulation
To solve the coupled Lindblad dynamics, we first rewrite the dynamics of the density operator in the superoperator formalism. We use the time-dependent density matrix renormalization group (TD-DMRG) method to evolve the density operators, which are propagated using the time-dependent variational principle (TDVP), implemented in Julia packages ITensors.jl and ITensorMPS.jl [58, 59]. As for the ordering of sites, we follow [7], in which we order the sites based on the magnitude of the dissipation. We set the cutoff threshold $\epsilon=10^{-12}$. After each TDVP step, we normalize the state to have a trace 1, i.e., $\operatorname{tr}(\hat{\rho})=1$. To evaluate any physical observable $\hat{O}$, we calculate the trace $\operatorname{tr}(\hat{O}\hat{\rho})$. We remark that to take traces of any operator $\hat{A}$, in the superoperator formalism, it means to calculate the inner product $\langle\langle I|A\rangle\rangle$, where $|I\rangle\rangle$ is the vectorization of the identity operator, and which can be explicitly constructed as a matrix product state.
@@ III TD-DMRG based simulation
IV Numerical experiments on the fermionic Anderson impurity model
@@
V Additional information on benchmarking BCF fitting
@@
In addition, here we show the fitting results in the time domain using $N=4$ modes in Fig. S2.
@@
VI Additional Information on calculating the absorption spectrum

### v2

@@ Abstract
Finally, we compute the absorption spectra using the coupled Lindblad pseudomode for a dimer model with three states, $\mathinner{|{g}\rangle},\mathinner{|{\epsilon_{1}}\rangle},$ and $\mathinner{|{\epsilon_{2}}\rangle}$, and $\hat{H}_{\text{S}}=\sum_{i=1}^{2}\epsilon_{i}\mathinner{|{\epsilon_{i}}\rangle}\!\mathinner{\langle{\epsilon_{i}}|}+J(\mathinner{|{\epsilon_{1}}\rangle}\!\mathinner{\langle{\epsilon_{2}}|}+\mathinner{|{\epsilon_{2}}\rangle}\!\mathinner{\langle{\epsilon_{1}}|})$. The environment is coupled independently to each excited state via the operators, $\hat{S}_{i}=\mathinner{|{\epsilon_{i}}\rangle}\!\mathinner{\langle{\epsilon_{i}}|}$, using the parameters of Ref. [27]. We consider two different spectral densities: $J_{0}(\omega)$, a broad spectrum, and $J_{1}(\omega)$, an additional sharp peak over $J_{0}(\omega)$ (details in SM [55, Section S7]). We compute the absorption spectrum, $S(\omega)=\omega\operatorname{Im}\left(\int_{0}^{\infty}\mathrm{i}C_{\hat{\mu}}(t)\mathrm{e}^{\mathrm{i}\omega t}\mathrm{d}t\right)$, derived from the dipole-dipole correlation function $C_{\hat{\mu}}(t)$ with $\hat{\mu}=\sum_{i}\mathinner{|{\epsilon_{i}}\rangle}\!\mathinner{\langle{g}|}+\mathinner{|{g}\rangle}\!\mathinner{\langle{\epsilon_{i}}|}$ and an initial state $\hat{\rho}_{\text{S}}(0)=\mathinner{|{g}\rangle}\!\mathinner{\langle{g}|}$. The result is shown in Fig. 3 compared to results from the Lorentzian pseudomode. Notably, the absorption spectrum exhibits a sharp peak in the negative frequency region at zero temperature, which becomes narrower and sharper as $N$ increases, highlighting the convergence behavior of our approach. Accurately capturing the presence of sharp resonances and broad features in the spectrum requires a faithful reconstruction of the spectral density, which our method achieves effectively. In contrast, the Lorentzian pseudomode fails to reproduce the broad component and misplaces several spectral peaks.
@@
Conclusions and outlook.– We argue that the coupled Lindblad pseudomode framework possesses all the desirable features of a well-designed pseudomode theory. The dynamics can be realized as a quantum channel, making them inherently stable. In both theory and practice, only a small number of pseudomodes is needed to achieve accurate results with $\text{polylog}(T/\varepsilon)$ scaling. Moreover, the pseudomodes can be constructed via a robust algorithm. We illustrate the method using the spin-boson model as an example, but it is equally applicable to fermionic environments. We anticipate that with the improved understanding and robustness provided by the techniques developed in this work, the coupled Lindblad framework will prove broadly useful across a wide range of open quantum system applications. These include the simulation of ultra-strong coupling regimes [56], the use of pseudomodes as an impurity solver in dynamical mean-field theory [57, 58, 17], and the simulation of condensed-phase chemical dynamics on quantum platforms such as trapped-ion devices [15, 16].
@@
(Dated: )
S1 Realization-based coupled Lindblad modes construction
In this section, we describe a new procedure for obtaining coupled Lindblad parameters directly from the BCF in the frequency domain, motivated by the realization problem in control theory [54, 42]. It is worth noting that the fitting of $C^{\text{q}}(t)$ in Eq. 5 is often performed in the time domain, using signal processing algorithms such as ESPRIT [63, 25] and Prony algorithm [64]. However, if the bath information is provided in the frequency domain rather than in the time domain, a Fourier transform of the spectral density is required, which can introduce additional approximation errors, particularly when $J(\omega)$ is only available on a discrete frequency grid with limited accuracy.
The realization problem aims to fit a given scalar function $f$ in the form $f(\omega)=l^{\dagger}(K-\omega I)^{-1}r$. The algorithm is highly robust and only requires applying a singular value decomposition (SVD) to the Loewner matrix [42] constructed from the sampled data $f(\omega)$.
$\widetilde{C}(\omega)\approx\mathrm{Im}(g^{\dagger}(K-\omega I)^{-1}g),\quad K=H-\mathrm{i}\Gamma.$ (S1)
S2 Coupled Lindblad pseudomodes theory for multi-site cases
@@
Therefore, the coupled Lindblad dynamics takes the following form:
[…]
@@
S3 Coupled Lindblad pseudomode theory for fermionic systems
@@
$\hat{H}_{\text{SB}}=\sum_{i=1}^{n}\int\mathrm{d}\omega f_{i}(\omega)\hat{a}_{i}^{\dagger}\hat{c}_{\omega}+\text{h.c.},$ (S5)
@@
Here $J(\omega)=f_{i}(\omega)\overline{f_{j}(\omega)}$ is the bath spectral density, $\mu$ is the chemical potential, $\beta$ is the inverse temperature, and $f_{\text{FD}}^{\mu,\beta}(\omega)=\frac{1}{1+\mathrm{e}^{\beta(\omega-\mu)}}$ is the Fermi-Dirac function. To apply the pseudomode theory for fermionic cases, the key is to account for both lesser and greater real-time hybridization functions arising in the problem. This is similar to what is done in quasi-Lindblad theories [25, Sec II. D]. We conduct fitting for both $\Delta^{<}(t)$ and $\Delta^{>}(t)$ as follows:
@@
$\hat{H}_{\text{A}_{1}}=\sum_{k,l=1}^{N_{1}}H^{<}_{kl}\hat{c}_{k}^{\dagger}\hat{c}_{l},\quad\hat{H}_{\text{SA}_{1}}=\sum_{i=1}^{n}\sum_{k=1}^{N_{1}}(\textbf{g}^{<})_{ki}\hat{c}_{k}^{\dagger}\hat{a}_{i}+\text{h.c.},$
$\hat{H}_{\text{A}_{2}}=\sum_{k,l=1}^{N_{2}}H^{>}_{kl}\hat{d}_{k}^{\dagger}\hat{d}_{l},\quad\hat{H}_{\text{SA}_{2}}=\sum_{i=1}^{n}\sum_{k=1}^{N_{2}}(\textbf{g}^{>})_{ki}\hat{d}_{k}^{\dagger}\hat{a}_{i}+\text{h.c.},$
@@
S4 TD-DMRG based simulation
[…]
@@
S5 Numerical experiments on the fermionic Anderson impurity model
@@
S6 Additional information on benchmarking BCF fitting
@@
In addition, here we show the fitting results in the time domain for the Ohmic spectral density using $N=4$ modes in Fig. S2.
@@
To further demonstrate the expressibility of the coupled Lindblad modes, here we also show the fitting results for sub-Ohmic densities $J(\omega)=\omega^{\alpha}\mathrm{e}^{-\omega/\omega_{c}}$ with $\alpha=1/2$ and $\omega_{c}=1.0$. Sub-Ohmic densities are known to be more challenging to simulate compared to Ohmic and super-Ohmic densitites due to its singularity at $\omega=0$, and exhibits very rich dynamical phases (see [65] for example). Our results in Fig. S3 exhibits the same $O(\mathrm{poly}\log(T/\varepsilon))$ scaling as in other examples, and achieve very accurate fitting with only $N=4$ bath modes.
Figure S3: Results for sub-Ohmic density $J(\omega)=\omega^{1/2}\mathrm{e}^{-\omega}$ using coupled Lindblad modes. (a) Number of coupled modes NN v.s. maximum simulation time $T$ for reaching $\varepsilon=10^{-5}$ accuracy. (b) Error v.s. number of coupled modes, for fixed maximum simulation time $T=10$. (c) Fitting results of $C(t)$, with $N=4$ coupled modes.
S7 Additional Information on calculating the absorption spectrum